# Advances in Embodied Navigation Using Large Language Models: A Survey

### v4

@@ 3.3.1 LLM-based Model
SQA3D is underpinned by an extensive dataset derived from 650 scenes in ScanNet, comprising 6.8k distinct scenarios, 20.4k textual descriptions, and 33.4k multifaceted reasoning questions. The scope and diversity of the dataset enable the exploration of complex questions that demand a nuanced understanding and facilitate the aggregation of a substantial data corpus.
To address SQA3D’s inherent challenges, researchers employed transformer-based vision-language models such as Scan-QA, which generates answers by associating 3D scans with questions. In addition, auxiliary tasks have been considered during the training phase to amplify model performance. Specialized models attuned to various 3D scene contexts, including egocentric video clips and bird-eye view images, have also been deployed. Notably, researchers have explored the potential of recent language and vision models, including GPT-3 and Unified QA, to tackle the SQA3D task in a zero-shot setting.
Experimental analyses involving state-of-the-art multi-modal reasoning models, such as ScanQA, were conducted to evaluate their efficacy on the SQA3D task against established benchmarks. The outcomes clearly establish SQA3D as a demanding task that necessitates complex reasoning about the agent’s environment for accurate question answering.
4)L3MVN [59] introduces a pioneering module framework that capitalizes on large language models to enhance visual target navigation. The framework aims to solve the challenge of autonomous exploration in unfamiliar settings, where a robot, provided with an object’s name, must proficiently navigate a 3D space to locate that object.
The architecture comprises two principal modules: a language module and a navigation module. The former handles natural language instructions, generating a semantic map embedded with general physical world knowledge. The latter employs this semantic map to guide robotic exploration, deducing the semantic pertinence of visible frontiers and opting for the most cost-efficient maneuvers.
L3MVN’s cardinal innovation resides in its utilization of large language models, specifically GPT-3, to imbue the system with common-sense reasoning for object location tasks. GPT-3, a leading-edge language model, undergoes fine-tuning on a limited set of task-specific data to produce a semantic map replete with object-searching-relevant information.
The theoretical underpinning of L3MVN is anchored in the notion that human adeptness in intricate and dynamic environments stems partially from inherent comprehension of the physical world. This common-sense knowledge can be transferred to robots via large language models, enhancing their operational efficiency in unknown terrains.
Efficacy evaluations of L3MVN were conducted on two comprehensive photorealistic 3D scene datasets, namely Gibson and HM3D. Experimental outcomes indicate substantial enhancements in both the success rate and efficiency of visual target navigation, while obviating the need for expansive learning processes. Additional ablation studies corroborate the efficacy of our language models and cost-utility exploration in facilitating more efficient navigation.
Figure 3: The second type utilizes LLMs to analyze incoming visual or textual data to extract goal-relevant information, upon which exploration policies subsequently generate appropriate actions to guide agent movement. LLMs, by acquiring information from text and other visual models processed through images, perform semantic understanding rather than planning. They extract key information sequences such as targets and locations, and hand them over to exploration algorithms. The exploration algorithms generate actions to guide agents in navigation to complete tasks.
5)ESC [47] (Exploration with Soft Commonsense Constrai-nts for Zero-shot Object Navigation) presents a groundbreaking approach to zero-shot object navigation by leveraging commonsense knowledge from pre-trained models for open-world navigation. Zhou et al. assert that effective object navigation hinges on two essential faculties: (1) Semantic scene understanding, crucial for recognizing objects and rooms, and (2) Commonsense reasoning, needed for logical inferences about probable locations of target objects based on general knowledge.
Previous zero-shot object navigation methods often fall short in commonsense reasoning. They either necessitate training on distinct goal-oriented navigation tasks and scenarios or depend on basic exploration heuristics. These approaches underperform when encountering novel objects or settings. Unlike these methods, ESC employs pre-trained visual and language models for open-world, hint-based grounding via the GLIP model [60], enabling scene comprehension and object grounding. Large-scale image-text pre-training further facilitates generalization to new objects.
@@
By setting a new technical work and pioneering the use of pre-trained commonsense knowledge in LLMs for object navigation, ESC paves the way for future research. This could include extracting intricate commonsense relationships, such as spatial connections between rooms, from LLMs. Additionally, future work might explore relaxing the zero-shot stipulations to permit limited fine-tuning in the learning of frontier selection strategies.
6)NavGPT [54] is an instruction-following visual navigation agent founded on Language Learning Models (LLMs). It aims to explore the reasoning capabilities of GPT models in intricate, embodied contexts via zero-shot sequential action prediction. Zhou et al. conducted extensive experiments to show that NavGPT excels in advanced navigation planning, including the dissection of instructions into sub-goals, the incorporation of common-sense knowledge pertinent to navigation tasks, landmark identification in observed scenes, navigation progress tracking, and adjustments to plans based on unexpected developments. The study also revealed LLMs’ aptitude for generating precise navigation instructions from observations and actions, as well as mapping accurate top-down metric trajectories from navigation history.Despite its robust capabilities, NavGPT’s performance in zero-shot Room-to-Room (R2R) tasks does not yet rival that of specialized models. To address this, they advocate the use of LLMs with multi-modal inputs as visual navigation agents and suggest applying LLMs’ explicit reasoning to learning-based models to boost performance.
NavGPT was introduced as a fully automated, LLM-based system designed for language-guided visual navigation. The system accommodates multi-modal inputs, unrestricted language guidance, interactions with open-world settings, and maintains a navigation history. Utilizing GPT-4 for intricate planning, NavGPT demonstrates proficiency in sequential action prediction. It excels in subdividing instructions into sub-goals, integrating navigation-relevant common-sense knowledge, recognizing landmarks in observed settings, tracking ongoing navigation progress, and making plan adjustments based on unexpected events. Furthermore, NavGPT demonstrates spatial and historical awareness by generating trajectory instructions and plotting navigated viewpoints in an overhead view.
@@
7)VELMA [62] is an LLM-based embodied intelligence agent designed for urban VLN in Street View settings. The agent navigates based on human-generated instructions, which include landmark references and directional cues. Schumann et al. developed VELMA to use verbal cues, contextualized by visual observations and trajectories, to determine subsequent actions. The system identifies landmarks from human-authored navigation instructions and employs CLIP to assess their visibility in the current panoramic view, achieving a linguistic representation of visual information.
LLMs have been increasingly utilized as reasoning engines for embodied agents in diverse applications, such as domestic robotics, video games, and indoor navigation. However, the challenge of interfacing optimally with interactive visual environments persists. This study addresses the issue through verbalization embodiment. Here, the agent’s task, trajectory, and visual observations are verbalized in natural language, allowing LLMs to comprehend navigation instructions and execute corresponding actions.
The paper introduces a landmark scorer to assess landmark visibility within panoramic images. This scorer calculates similarity metrics between textual descriptions of landmarks and their visual representations, using CLIP models. Each landmark receives a normalized score based on visual similarity. If this score exceeds a predefined threshold, the landmark is deemed visible.
The landmark classification is unsupervised, given the absence of ground-truth labels for visibility. The scorer’s threshold is its only tunable parameter. The agent also assesses views to its current left and right orientations. Each directional view is linked to specific textual information. Visible landmarks, along with their associated directional texts, are fed into a language expresser, which then generates verbal descriptions of the environment.
These verbal observations are of two kinds: street intersections and landmarks visible within the current view. These strings encode information about the number of outgoing edges at the current node, the names of visible landmarks, and their associated directional texts.
The study evaluated the performance of the proposed LLM agent in both fine-tuning and contextual learning scenarios. A GPT-3-based landmark extractor was used for a single run and its extracted landmarks were employed in all subsequent experiments. For the landmark scorer, the CLIP-ViT-bigG-14-laion2B-39B-b160k model was chosen as the CLIP model.
@@
8) $A^{2}$av [63] is an action-aware vision-and-language navigation method that consists of an instruction parser and an action-aware navigation strategy. The instruction parser, which employs large language models such as GPT-3, decomposes complex navigation instructions into a sequence of object-specific navigation sub-tasks that relate to particular actions. These sub-tasks necessitate that the agent identify objects and navigate to target locations based on associated action requirements. Despite advancements in VLN, current approaches have failed to sufficiently account for diverse action requirements within instructions, such as ”continue beyond” and ”exit”, thus ignoring the agent-object relationship in the scene.
To tackle this problem, the agent must accurately identify and execute the action requirements associated with each milestone. However, the complexity of grammatical structures and the diversity of action expressions in instructions make understanding these requirements challenging. Additionally, the development of navigation strategies capable of meeting these action requirements without annotated path instructions remains an open question.
$A^{2}$av employs the reasoning capabilities of large language models to parse instructions into action-specific object navigation sub-tasks. In these sub-tasks, the agent identifies landmarks and navigates based on pertinent action requirements. Addressing the limitations of existing zero-shot navigation agents, Chen et al. identified five fundamental action requirements and developed distinct navigators for each. This approach, which obviates the need for manual path instruction annotations, yields more accurate and interpretable navigation outcomes. The system utilizes GPT-3 to parse all landmarks and their associated action requirements, effectively decomposing the instruction into a series of sub-tasks specific to actions.
To execute these sub-tasks sequentially, $A^{2}$av adopts a zero-shot learning approach, comprising five action-specific navigators. While parsing instructions, GPT-3 predicts the ”ACTION”, which typically aligns with one of the predefined ”ACTION” for a sub-task, allowing direct mapping to the corresponding sub-task type. Chen et al. encode all predicted and predefined ”ACTION” using BERT [64] and select the sub-task with the highest cosine similarity score as the predicted sub-task.
For action-aware navigation, Chen et al. fine-tuned ZSON [65] to individually train five distinct action-specific navigators. In each specific action dataset, the positions of landmarks and paths reveal features pertinent to the respective action requirement. In any given task, the current sub-task type is first identified, and the corresponding action-specific navigator is selected to predict low-level actions. The process iterates until either all sub-tasks are completed or the maximum step limit for the episode is reached.
In summary, $A^{2}$av effectively addresses action requirements within VLN tasks by breaking them down into action-specific object navigation sub-tasks. It achieves superior zero-shot VLN performance on the RxR and R2R datasets and outperforms state-of-the-art supervised methods on RxR. Additionally, $A^{2}$av excels in navigating according to instructions with specific action requirements, underscoring its potential in scenarios requiring human-agent interactions.
Table 2: Performance Comparison (Part 1)
Benchmark Habitat RoboTHOR Gibson
SPL SR SPL SR SPL SR
CoW 6.3% - 10% 16.3% - -
ZSON - - - - 28% 36.9%
L3MVN - - - - 37.7% 76.1%
VLMAP 6% 2.5% - - - -
ESC - - 22.2% 38.1% - -
9)MiC [55] (March in Chat) is an environment-aware instruction planner that employs an LLM for dynamic dialogues, specifically designed for the REVERIE dataset. The architecture of MiC is trifurcated into three modules: Generalized Object-and-Scene-Oriented Dynamic Planning (GOSP), Step-by-step Object-and-Scene-Oriented Dynamic Planning (SODP), and Ro-om-and-Object Aware Scene Perceiver (ROASP).
@@
In summary, MiC provides concise high-level instructions to VLN agents, enabling REVERIE agents to dynamically interact with the LLM to formulate plans for upcoming steps. This dynamic planning methodology posits a promising avenue for future research and development.
10)SayNav [53] is a groundbreaking framework designed to harness the common-sense knowledge encapsulated in Large Language Models (LLMs) for generalizing intricate navigation tasks in expansive and unfamiliar environments. SayNav employs a unique grounding mechanism that incrementally constructs a 3D scene graph of the explored territory, subsequently feeding this information to LLMs. This results in the generation of high-level navigation plans that are context-aware and practically implementable, executed thereafter by a pretrained low-level planner as a series of short-distance point-goal navigation sub-tasks.
@@
The low-level planner is tasked with generating brief motion control commands. It operates on RGBD images (320x240 resolution) and the agent’s pose data, outputting basic navigational commands in alignment with standard PointNav configurations. To unify the operations of the high-level and low-level planners, SayNav treats each pseudo-code instruction from the LLM as a short-distance point-goal navigation sub-task. The 3D coordinates of the objects specified in each planned step determine the target points for these sub-tasks.
@@
SayNav distinguishes itself by its ability to dynamically generate navigation instructions and iteratively refine its future steps based on newly collected data. Performance metrics on multi-object navigation tasks demonstrate its substantial advantage over Oracle-based PointNav methods, confirming its proficiency in dynamically planning and executing object-finding tasks in large, unfamiliar environments. Moreover, SayNav has proven its efficacy in generalizing from simulated to real-world conditions. In conclusion, SayNav emphasizes the utilization of human-like reasoning and advanced semantic understanding for achieving efficient and adaptive navigation, especially in complex or uncharted settings.
Table 3: Performance Comparison (Part 2)
Benchmark HM3D MP3D REVERIE R2R
SPL SR SPL SR SPL SR SPL SR
ZSON 12.6% 25.5% 4.8% 15.3% - - - -
CLIP-Nav - - - - 4.56% 15.2% - -
L3MVN - - 23.1% 50.4% - - - -
OVRL 29% 64% - - - - - -
ESC 22% 38.5% 13.7% 28.6% - - - -
NavGPT - - - - - - 29% 34%
A2Nav - - - - - - 11% 22.6%
MIC - - - - 41.97% 55.74% - -
SayNav 34% 60.32% - - - - - -
@@
LLMs have demonstrated significant contributions to navigation tasks, primarily through their enhanced capabilities in environmental perception and decision-making. These models integrate advanced language and image processing technologies to provide sophisticated environmental perception and decision support, crucial for tasks requiring a deep understanding of the environment and quick, accurate decision-making. Unlike traditional navigation models that are limited to predefined rules and instructions and offer limited interaction modes, LLMs excel by enabling more dynamic and flexible route planning. Researchers are continuously exploring various applications of LLMs in embodied navigation systems to further enhance their effectiveness in navigation tasks. These applications extend beyond traditional navigation functionalities to include augmented reality perception, multimodal interaction, procedural generation, and automated decision support.
@@
Given the rapid development of LLMs, several studies have explored methods and practices to enhance traditional navigation tasks using these technologies. LLMs can serve as decision-making components in complex scenarios within autonomous driving systems, by designing cognitive paths and developing algorithms to translate LLM decisions into executable driving commands. This approach not only improves performance in single-vehicle tasks but also handles complex driving behaviors and multi-vehicle coordination, thanks to the common-sense reasoning capabilities of LLMs [68].
@@ 3.3.2 Other Models
1)CoW [74] (CLIP on Wheels) by Gadre et al. [74] is an innovative approach that adapts zero-shot visual models like CLIP to embodied AI tasks, particularly object navigation. The framework involves an agent identifying a target object in unfamiliar environments, defined through text. The key concept is dividing the task into zero-shot object localization and exploration.
@@
2)ZSON [65] introduces an innovative methodology for instructing virtual robots to navigate unfamiliar terrains and identify objects without pre-existing rewards or demonstrations. Distinct from conventional ObjectNav techniques, ZSON exploits image-goal navigation (ImageNav) to transcode goal images into a multimodal, semantic embedding space. This allows for the scalable training of semantic-goal navigation (SemanticNav) agents in unannotated 3D settings. The underpinning theory of ZSON hinges on the principle of semantic similarity. By encoding goal images as semantic embeddings, the approach enables agents to navigate toward objects predicated on their semantic likeness to the goal image. This strategy assumes that objects bearing semantic resemblance to the goal image are likely located in similar spatial contexts.
To conclude, the approach put forth by the researchers yielded commendable success metrics across the three ObjectNav datasets, constituting a significant divergence from traditional ObjectNav strategies. The model is scalable, zero-shot, and well-suited for application in unannotated 3D worlds, establishing it as a viable avenue for open-world ObjectNav implementations.
3)VLMAP [76] presents an innovative spatial map representation that integrates pretrained visual-language features with a 3D environmental reconstruction, facilitating natural language map indexing without the need for extra labeled data. The central innovation is its capacity to intrinsically combine visual-language attributes with the 3D environmental blueprint, empowering robots to execute spatially precise navigation via natural language directives.
VLMAP advances the fields of computer vision, natural language processing, and robotics by unifying visual-language features with 3D reconstructions for precise, natural language-guided robotic navigation. The incorporation of visual-language features leverages recent progress in deep learning, while the 3D reconstruction component builds on established principles of 3D point cloud processing. The natural language capabilities are supported by an expanding corpus of research in natural language processing and human-robot interaction.
Experimental evaluations of VLMAP substantiate its efficacy in enabling robots to navigate spatially precise paths through natural language commands. Tests in a simulated environment affirm the system’s capability to convert natural language directives into a sequence of open-vocabulary navigation objectives that can be directly pinpointed on the map. Further, the adaptability of VLMaps for use among robots with disparate configurations to generate real-time obstacle maps is also demonstrated.
Table 4: Comparison of Works This table provides a comparison of different Works, the design structure, and their respective applications. It’s worth noting that each wok uses different evaluation criteria, which are not listed here. For specific details on each work, please refer to the citations at the end of the article.
@@
4)OVRL [77] introduces a neural network architecture specifically aimed at visual navigation challenges. Notably, the architecture is composed of task-agnostic elements such as ViTs, convolutions, and LSTMs, obviating the need for task-specific modules. The standout feature of OVRL-V2 is a novel compression layer designed to preserve spatial details in visual navigation assignments. This layer compresses high-dimensional image attributes from the ViT model into low-dimensional features, maintaining the essential spatial context.
@@
Experimental validations affirm OVRL-V2’s state-of-the-art performance in both IMAGENAV and OBJECTNAV tasks, registering success rates of 82.0% and 64.0%, respectively. These results signify marked advancements over preceding methodologies, corroborating the efficacy of both the proposed architecture and its associated training techniques.
@@
3.3.3 Comparison
$\text{SPL}=\frac{1}{N}\sum_{i=1}^{N}S_{i}\frac{L_{i}}{\max(P_{i},L_{i})}$
We have compared the works mentioned above, evaluating aspects such as the large models employed, application domains, and additional features. These comparisons are summarized in Table 4.At the same time, we have divided the aforementioned works into two categories: Planning, as illustrated in fig.4, and Semantic Understanding, as shown in fig.5. Corresponding examples are provided for each category. Furthermore, we have analyzed the performance of these works across various datasets, the details of which are provided in Tables 2 and 3. It is important to clarify that comprehensive evaluation was unfeasible for some works due to the unavailability of public code. Consequently, the results are based solely on performance metrics reported in the corresponding publications. For an in-depth understanding of each work, readers are directed to the cited literature.
Table 5: Evaluation Metrics The metrics of vision-and-language navigation [7].

### v5

@@ III-C1 LLM-based Model
d) ESC [180] (Exploration with Soft Commonsense Constrai-nts for Zero-shot Object Navigation) presents a approach to zero-shot object navigation by leveraging commonsense knowledge from pre-trained models for open-world navigation. Zhou et al. assert that effective object navigation hinges on two essential faculties: (1) Semantic scene understanding, crucial for recognizing objects and rooms, and (2) Commonsense reasoning, needed for logical inferences about probable locations of target objects based on general knowledge.
Unlike previous methods, ESC employs pre-trained visual and language models for open-world, hint-based grounding via the GLIP model [68], enabling scene comprehension and object grounding. Large-scale image-text pre-training further facilitates generalization to new objects.
@@
e) NavGPT [179] aims to explore the reasoning capabilities of GPT models in intricate, embodied contexts via zero-shot sequential action prediction. Zhou et al. conducted extensive experiments to show that NavGPT excels in advanced navigation planning, including the dissection of instructions into sub-goals, the incorporation of common-sense knowledge pertinent to navigation tasks, landmark identification in observed scenes, navigation progress tracking, and adjustments to plans based on unexpected developments. The study also revealed LLMs’ aptitude for generating precise navigation instructions from observations and actions, as well as mapping accurate top-down metric trajectories from navigation history.
The system accommodates multi-modal inputs, unrestricted language guidance, interactions with open-world settings, and maintains a navigation history. Utilizing GPT-4 for intricate planning, NavGPT demonstrates proficiency in sequential action prediction. It excels in subdividing instructions into sub-goals, integrating navigation-relevant common-sense knowledge, recognizing landmarks in observed settings, tracking ongoing navigation progress, and making plan adjustments based on unexpected events. Furthermore, NavGPT demonstrates spatial and historical awareness by generating trajectory instructions and plotting navigated viewpoints in an overhead view.
@@
f) VELMA [107] is designed for urban VLN in Street View settings. The agent navigates based on human-generated instructions, which include landmark references and directional cues. VELMA used verbal cues, contextualized by visual observations and trajectories, to determine subsequent actions. The system identifies landmarks from human-authored navigation instructions and employs CLIP to assess their visibility in the current panoramic view, achieving a linguistic representation of visual information.
VELMA introduces a landmark scorer to assess landmark visibility within panoramic images. This scorer calculates similarity metrics between textual descriptions of landmarks and their visual representations, using CLIP models. Each landmark receives a normalized score based on visual similarity. If this score exceeds a predefined threshold, the landmark is deemed visible.
The landmark classification is unsupervised, given the absence of ground-truth labels for visibility. The scorer’s threshold is its only tunable parameter. The agent also assesses views to its current left and right orientations. Each directional view is linked to specific textual information. Visible landmarks, along with their associated directional texts, are fed into a language expresser, which then generates verbal descriptions of the environment.These verbal observations are of two kinds: street intersections and landmarks visible within the current view. These strings encode information about the number of outgoing edges at the current node, the names of visible landmarks, and their associated directional texts.
@@
f) MiC [96] (March in Chat) is an environment-aware instruction planner that employs an LLM for dynamic dialogues, specifically designed for the REVERIE dataset. The architecture of MiC is trifurcated into three modules: Generalized Object-and-Scene-Oriented Dynamic Planning (GOSP), Step-by-step Object-and-Scene-Oriented Dynamic Planning (SODP), and Ro-om-and-Object Aware Scene Perceiver (ROASP).
@@
g) SayNav [101] is a groundbreaking framework designed to harness the common-sense knowledge encapsulated in Large Language Models (LLMs) for generalizing intricate navigation tasks in expansive and unfamiliar environments. SayNav employs a unique grounding mechanism that incrementally constructs a 3D scene graph of the explored territory, subsequently feeding this information to LLMs. This results in the generation of high-level navigation plans that are context-aware and practically implementable, executed thereafter by a pretrained low-level planner as a series of short-distance point-goal navigation sub-tasks.
@@
The low-level planner is tasked with generating brief motion control commands. It operates on RGBD images (320x240 resolution) and the agent’s pose data, outputting basic navigational commands in alignment with standard PointNav configurations[63]. To unify the operations of the high-level and low-level planners, SayNav treats each pseudo-code instruction from the LLM as a short-distance point-goal navigation sub-task. The 3D coordinates of the objects specified in each planned step determine the target points for these sub-tasks.
@@
SayNav distinguishes itself by its ability to dynamically generate navigation instructions and iteratively refine its future steps based on newly collected data. Performance metrics on multi-object navigation tasks demonstrate its substantial advantage over Oracle-based PointNav methods, confirming its proficiency in dynamically planning and executing object-finding tasks in large, unfamiliar environments. Moreover, SayNav has proven its efficacy in generalizing from simulated to real-world conditions. In conclusion, SayNav emphasizes the utilization of human-like reasoning and advanced semantic understanding for achieving efficient and adaptive navigation, especially in complex or uncharted settings. h) ETPNav [5] combines topological and metric map construction to balance long-term planning and short-term decision-making. The system processes panoramic RGB images to generate feature vectors and uses a pre-trained ViT to extract image features, while also incorporating depth information to build a metric map for real-time navigation decisions. The topological map records all observed nodes, supporting long-term planning, and represents each node’s spatial position and environmental information through contextual view embeddings. The metric map is locally constructed around the current node, combining grid features from nearby visited nodes to accurately represent the spatial details of the environment for immediate navigation.
Additionally, the framework employs a cross-modal long-term and short-term Transformer model for both long-term planning and real-time navigation. The long-term Transformer handles long-term planning tasks by combining node embeddings and topological map encoding, while the short-term Transformer optimizes real-time decisions by integrating metric map data and current node information. This multi-modal fusion model enhances the agent’s navigation ability in complex environments, enabling it to effectively perform tasks that require both visual and language understanding.
@@
LLMs have demonstrated significant contributions to navigation tasks, primarily through their enhanced capabilities in environmental perception and decision-making. These models integrate advanced language and image processing technologies to provide sophisticated environmental perception and decision support, crucial for tasks requiring a deep understanding of the environment and quick, accurate decision-making. Unlike traditional navigation models that are limited to predefined rules and instructions and offer limited interaction modes, LLMs excel by enabling more dynamic and flexible route planning. Researchers are continuously exploring various applications of LLMs in embodied navigation systems to further enhance their effectiveness in navigation tasks [23]. These applications extend beyond traditional navigation functionalities to include augmented reality perception, multimodal interaction, procedural generation, and automated decision support.
@@ III-C2 Other Models
a) CoW [43] (CLIP on Wheels) by Gadre et al. [43] is an innovative approach that adapts zero-shot visual models like CLIP to embodied AI tasks, particularly object navigation. The framework involves an agent identifying a target object in unfamiliar environments, defined through text. The key concept is dividing the task into zero-shot object localization and exploration.
@@
TABLE II: Comparison of Works This table provides a comparison of different Works, the design structure, and their respective applications. It’s worth noting that each wok uses different evaluation criteria, which are not listed here. For specific details on each work, please refer to the citations at the end of the article.
@@
b) ZSON [83] introduces an innovative methodology for instructing virtual robots to navigate unfamiliar terrains and identify objects without pre-existing rewards or demonstrations. Distinct from conventional ObjectNav techniques, ZSON exploits image-goal navigation (ImageNav) to transcode goal images into a multimodal, semantic embedding space. This allows for the scalable training of semantic-goal navigation (SemanticNav) agents in unannotated 3D settings. The underpinning theory of ZSON hinges on the principle of semantic similarity. By encoding goal images as semantic embeddings, the approach enables agents to navigate toward objects predicated on their semantic likeness to the goal image. This strategy assumes that objects bearing semantic resemblance to the goal image are likely located in similar spatial contexts.
c) VLMAP [51] presents an innovative spatial map representation that integrates pretrained visual-language features with a 3D environmental reconstruction, facilitating natural language map indexing without the need for extra labeled data. The central innovation is its capacity to intrinsically combine visual-language attributes with the 3D environmental blueprint, empowering robots to execute spatially precise navigation via natural language directives.
VLMAP advances the fields of computer vision, natural language processing, and robotics by unifying visual-language features with 3D reconstructions for precise, natural language-guided robotic navigation. The incorporation of visual-language features leverages recent progress in deep learning, while the 3D reconstruction component builds on established principles of 3D point cloud processing. The natural language capabilities are supported by an expanding corpus of research in natural language processing and human-robot interaction.Experimental evaluations of VLMAP substantiate its efficacy in enabling robots to navigate spatially precise paths through natural language commands. Tests in a simulated environment affirm the system’s capability to convert natural language directives into a sequence of open-vocabulary navigation objectives that can be directly pinpointed on the map. Further, the adaptability of VLMaps for use among robots with disparate configurations to generate real-time obstacle maps is also demonstrated.
d) OVRL [152] introduces a neural network architecture specifically aimed at visual navigation challenges. Notably, the architecture is composed of task-agnostic elements such as ViTs, convolutions, and LSTMs, obviating the need for task-specific modules. The standout feature of OVRL-V2 is a novel compression layer designed to preserve spatial details in visual navigation assignments. This layer compresses high-dimensional image attributes from the ViT model into low-dimensional features, maintaining the essential spatial context.
@@
III-C3 Comparison
TABLE III: Evaluation Metrics The metrics of vision-and-language navigation [115].